OpenStudy (anonymous):

Help. Will give medal and become fan for the best answer! :) Solve –4the square root of x plus two. + 3 = –9

OpenStudy (solomonzelman):

\[\sqrt[-4]{x+2}+3=-9\] this?

OpenStudy (anonymous):

Yep!

OpenStudy (solomonzelman):

\[\huge\color{blue}{ \sqrt[-4]{x+2}+3=-9 } \]\[\huge\color{blue}{ \sqrt[-4]{x+2}+3\color{red}{-3 } =-9 } \color{red}{-3 } \]\[\huge\color{blue}{ \sqrt[-4]{x+2}=-12 }\]

OpenStudy (solomonzelman):

\[\huge\color{blue}{ \frac{1 }{\sqrt[4]{x+2}}=-12 }\]

OpenStudy (solomonzelman):

\[\huge\color{blue}{ \frac{\sqrt[4]{1} }{\sqrt[4]{x+2}}=-12 }\]

OpenStudy (solomonzelman):

\[\huge\color{blue}{ \sqrt[4]{\frac{1}{x+2}}=-12 }\] \[\huge\color{blue}{ \frac{1}{x+2}=-12 ^4}\]

OpenStudy (solomonzelman):

I am not 100% sure, but think this is the way to do it.

OpenStudy (anonymous):

x = –1 x = 7 x = 31 x = 46 So which answer would it be?

OpenStudy (solomonzelman):

ooops, I was wrong then.

OpenStudy (anonymous):

If you had to guess which would you pick? Its my last question for Algebra 2 and I just need to finish lol

OpenStudy (solomonzelman):

plug in the choices.

OpenStudy (anonymous):

Alright, thank for your help!

OpenStudy (solomonzelman):

Yw

OpenStudy (anonymous):

there is no such thing as \[\sqrt[-4]{x+2}+3=-9\]

OpenStudy (anonymous):

maybe it is \[-4\sqrt{x+2}+3=-9\] in which case you can start with \[-4\sqrt{x+2}=-12\] then divide by \(-4\) to get \[\sqrt{x+2}=3\] square both sides giving \[x+2=9\] solve for \(x\)